OpenStudy (anonymous):

What is the inverse of the given relation? y = 7x2 – 3

OpenStudy (anonymous):

y+2=7x^2 so (y+2/7)=x^2 so x=sqr ((y+2/7)) then x replace to y then we have: y=sqr ((x+2/7))

OpenStudy (anonymous):

Oh whoops!! the 2 is an exponent! Totally my fault!!

OpenStudy (anonymous):

So, what is the inverse of 7x^2-3,

OpenStudy (anonymous):

after operation above: we have: y=sqr ((x+2/7))

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

Thank you so much!!!!!! And to find the inverse of a problem such as y=-4x^2-4 or something, I would do the opposite?